OpenStudy (tylermckinney16):

How many ways can 6 students be assigned to the first, second and third places in the lunch line?

jimthompson5910 (jim_thompson5910):

Think of three slots A,B,C There are 6 ways to pick someone to fill slot A After you pick one person, there are 5 people left. So there are 5 people for slot B Finally, there are 4 people for slot C 6*5*4 = 30*4 = 120 so there are 120 different ways to pick 3 people where order matters from a pool of 6 overall.

OpenStudy (tylermckinney16):

That the answer?

OpenStudy (tylermckinney16):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes there are 120 ways to fill those 3 slots (from a pool of 6 people)